Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

NSRA Centenary Postal Competition 2001

Distribution

Dr. Trevor West DUCAC Chairman
Cathy Doyle DUCAC Administrative Officer
Geraldine McAuley DUCAC Administrative Officer
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Helen Murphy DUCAC Sports Facilities Committee Secretary

Summary

Venue Postal Competition, shot in TCD, scored in England
Date October 2000–June 2001
Discipline 25yd Prone Rifle and 10m Air Rifle

DURC Attendees

Mark Dennehy  Ray Reilly
John Keeney  Ben Jones
Alan Smyth  Aisling Broderick

Scores Round One 25yd Prone Rifle

Ray Reilly Class A  395
John Keeney Class C  390
Alan Smyth Class D  385
Mark Dennehy Class B  381

10m Air Rifle

Ben Jones Class D  331
Aisling Broderick Class B  293

Round Two 25yd Prone Rifle

Ray Reilly Class A - Joint 55th  390
John Keeney Class C - Joint 25th  387
Alan Smyth Class D - Joint 21st  383

10m Air Rifle

Ben Jones Class D - 6th, continues to Final 334

Report

The NSRA is one of the largest and oldest existing association in the Target Shooting sport. This year marks it’s 100th anniversary (by contrast, DURC is only 40 this year). To celebrate, the NSRA have held a Centenary Postal Competition. Postal competitions differ from shoulder-to-shoulder competition in that the shooters do not travel to other ranges to shoot their cards, but shoot them on their home ranges and post the cards to the central scorer. DURC has entered these competitions on a regular basis for over twenty years - however, this competition has proven to be the largest we have ever entered and the most prestigous.

There were two rounds shot - one in March and the other in May this year. Ray Reilly, John Keeney and Alan Smyth successfully got to the second round for smallbore and Ben Jones made it through for Air Rifle. However, the standard of competition was extremely high, and none of the smallbore competitors made it to the final - Ray missing by 4 points in class A, John by 2 points in class C and Alan by 1 point in class D – all very, very close indeed, and a good acquittal of themselves, as they were competing against the shooters from all over england - several hundred in total.

Ben Jones, however, not only placed very high in his class (6th place in class D), but also qualified for the final in Bisley in early August. This is a remarkable achievement, and will be recorded in the sport’s history books as this event is a very prestigous one. We extend our congradulations to Ben, and hope he places highly in the finals and we will report on his progress at that point.

Misé le meas,

Mark Dennehy, DURC Secretary